Operations Report – Sept 4th

Off season  Operations Report 4 Sept 2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Hab water filter is black

Notes on non-nominal systems:  the water filter is black. I don’t know that replacing it now would do anything so we will need to clean the system before any use in Oct.

Generator (hours run): not run

Generator SOC currently 89% at 6 pm

Diesel Reading – 71%

Propane Reading – Shannon’s  20%.  Main tank 41 %

Ethanol Free Gasoline – 8 gallons.

Water (Axillary tank)  – 0 gallons

Water (static tank) – 400 gallons. 150 gallons at Shannon’s

Axillary to Static tank transfer– no

Gallons transferred: none

Water in GreenHab – n/a gallons

Water (loft) – Static to Loft Pump used – no

Water Meter: n/a

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Notes on rovers: all working rovers in town being used and well cared for.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

No

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: zero

Notes on ATVs: nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Yes. To town for drinking water.

General notes and comments:  Productive meeting with Ikea (SLC).  We worked on changes to the plans provided by Ikea in Sweden to improve function over form. The plan is to do the upgrades in three stages: lower deck and GreenHab first to be completed during the October work party. A new kitchen and science lab will be installed later in the season.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome: Dual split unit was installed in the interior but need the pad for the outside unit which did not get ordered. Once that is herewe will contact an Hvac company to add the refrigerant and then the Science Dome will have both ac and heat. This was installed mainly so we can keep the power system cool in the summer as the dome is pretty well insulated.

Summary of RAM operations: nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: continues to be unseasonably warm. I either have a cold or allergies.

Questions, concerns and requests to Mission Support:

Correction from last night’s report: power lost on June 27th, not July.